# CURSOR: Scalable Mixed-Order Hypergraph Matching with CUR Decomposition

### v2

@@ 1 Introduction
Figure 1: The comparison between the traditional ANN-based framework and CURSOR. Instead of calculating the whole tensor block (the light orange area in 0(a)) and extracting the highest compatibilities in each block (the blue cubes), CURSOR only calculates a small number of block fibers (the light orange area in 0(b)) and retains fewer elements in these fibers, effectively reducing computational costs for large-scale hypergraph matching. The method chooses the fibers based on the second-order matching result. CURSOR calculates fibers in all three tensor modes, and only one is shown in 0(b) for clarity.
@@ 3.1 Problem Setup
$\max_{\mathbf{x}}\quad$ $\mathbf{x}^{T}\mathbf{H}\mathbf{x}$ (2)
$s.t.\quad$ $\sum_{j\in\text{ind}_{i}}x_{j}=1,\forall i$
@@
$\max_{\mathbf{x}}\quad$ $\mathcal{H}\otimes_{1}\mathbf{x}\otimes_{2}\cdots\otimes_{k}\mathbf{x}$ (3)
$s.t.\quad$ $\sum_{j\in\text{ind}_{i}}x_{j}=1,\forall i$
@@ 4 Experiments
Table 1: Results on the synthetic dataset with ADGM and CURSOR.
@@ 4.2 Templates with Specific Shapes
Figure 2: Matching result with deformation 1(a) rotation with angle $[-30^{\circ},30^{\circ}]$, 1(b) scale on $x$-coordinate with scale factor $1.1^{\beta}$ where $\beta\in[-5,5]$, 1(c) adding noise with $\sigma/\sigma_{0}=[0,0.1]$, and 1(d) adding $n_{l}$ outliers where the outlier ratio$=n_{l}/n_{1}$.
@@ 4.3 CMU House and Hotel Dataset
Figure 4: The Cars and Motorbikes datasets with CURSOR and state-of-the-art hypergraph matching algorithms. 3(a) The number of non-zero compatibilities with ANN-based methods and CURSOR. The matching accuracy on the 3(b) Motorbikes and 3(c) Cars datasets.
@@ 6 PRL-based Method Derivation
Figure 6: The detailed hyperedge correspondences between one sampled hyperedge from the source (yellow triangle in 5(a)) and target images with ANN (5(b) and 5(c)) and CURSOR 5(d). The white triangles in the target images denote all the hyperedges compared with the source hyperedge, and the blue triangles are the hyperedges with the highest compatibilities ($r_{1}$ for ANN and $r$ for CURSOR). The green dashed triangles represent the matched hyperedges, and the red one represents the mismatch.
@@
We provide a more detailed analysis of the experiment results from Sec. 4.3 to demonstrate the superiority of CURSOR, as is shown in Fig. 6. Traditional ANN-based methods compute the compatibilities between the sampled source hyperedges (yellow triangle in Fig. 5(a)) and all the target ones (fully connected white triangles in Figs. 5(b)-5(c)). With the intermediate second-order result, CURSOR computes fewer compatibilities, as is shown in Fig. 5(d). To find the corresponding hyperedge (green dashed triangles in Fig. 6), a large amount (the hyperparameter $r_{1}=900$ in Fig. 5(b)) of the highest compatibilities (blue triangles in Figs. 5(b)-5(d)) should be selected for ANN. If we decrease $r_{1}$ to 50, some correct ones will be missed (red triangle in Fig. 5(c)). CURSOR can effectively find the hyperedges with the 25 highest compatibilities (green triangle in Fig. 5(d)). Compared to the traditional tensor generation methods, CURSOR can effectively increase the matching performance with less computational complexity.
@@ 8.1 CUR-based Pairwise Matching
Figure 7: 6(a) CUR-based second-order graph matching accuracy with various $c$. 6(b) The average accuracy using CURSOR with CUR-based pairwise matching result and CURSOR with randomly sampled indices (CURSOR-R).
@@ 8.2 CURSOR vs ANN-based Tensor Generation
Figure 8: Results on 100-vs-110 synthetic dataset comparing CURSOR with ANN, $\alpha=0.2$ for CURSOR. 7(a) The average matching accuracy using CURSOR and ANN with different $\alpha$. 7(b) The average iterations for convergence using CURSOR and ANN with different $\alpha$. 7(c) The decay $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}$ per iteration with different $\alpha$ using CURSOR and ANN. The noise level $\sigma=0.1$.
@@ 9.1 Parameter in Second-order Graph Matching
Figure 9: 8(a) The average hit rate over increasing $k$ on a 100-vs-110 synthetic dataset with $\sigma=0.02$ from the CUR-based second-order graph matching results. 8(b) The average accuracy on the 100-vs-110 synthetic dataset with $\sigma\in[0,0.2]$ and $c\in[10,100]$.
@@ 9.2 Parameters During Tensor Generation
Figure 10: Results on 100-vs-110 synthetic dataset with $\sigma\in[0,0.2]$ using CURSOR. 9(a) The matching accuracy with increasing $k$. 9(b) The matching accuracy with increasing $r$.
Figure 11: Results on 100-vs-110 synthetic dataset with PRL-based matching algorithm. 10(a) The average matching accuracy using CURSOR with different $\alpha$. 10(b) The average iterations for convergence using CURSOR with different $\alpha$. 10(c) The decay $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}$ per iteration with different $\alpha$ settings using CURSOR. The noise level $\sigma=0.1$.
@@ 9.3 Parameters of PRL-based Matching Algorithm
To study the sensitivity of $\alpha$, the matching accuracy and convergence speed using CURSOR were further analyzed. During the experiment, $c=100$, $k=10$, and $r$ was set as $100$. The stopping criteria of the PRL-based algorithm was assigned as $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}\leq 10^{-8}$. The matching accuracy with various $\alpha$ is shown in Fig. 10(a). With a reliable compatibility tensor, the PRL-based matching algorithm achieved almost the same performance regardless of $\alpha$. We further analyzed their convergence speeds, as shown in Figs. 10(b) and 10(c). As the noise level increased, more iterations were required to satisfy the stopping criteria, and a lower $\alpha$ achieved faster convergence. As discussed in the previous sections, the parameter $\alpha$ is a balanced factor between first and third-order compatibilities. Although the method converged faster with $\alpha=0$, the first-order compatibilities stabilized the matching process and increased the matching performance under some extreme circumstances.

### v4

@@ 1 Introduction
Figure 1: The comparison between the traditional ANN-based framework and CURSOR. Instead of calculating the whole tensor block (the light orange area in 1(a)) and extracting the highest compatibilities in each block (the blue cubes), CURSOR only calculates a small number of block fibers (the light orange area in 1(b)) and retains fewer elements in these fibers, effectively reducing computational costs for large-scale hypergraph matching. The method chooses the fibers based on the second-order matching result. CURSOR calculates fibers in all three tensor modes, and only one is shown in 1(b) for clarity.
@@ 3.1 Problem Setup
$\max_{\mathbf{x}}$ $\mathbf{x}^{T}\mathbf{H}\mathbf{x}$ (2)
$s.t.$ $\sum_{j\in\text{ind}_{i}}x_{j}=1,\forall i$
@@
$\max_{\mathbf{x}}$ $\mathcal{H}\otimes_{1}\mathbf{x}\otimes_{2}\cdots\otimes_{k}\mathbf{x}$ (3)
$s.t.$ $\sum_{j\in\text{ind}_{i}}x_{j}=1,\forall i$
@@ 4.2 Templates with Specific Shapes
Figure 2: Matching result with deformation 2(a) rotation with angle $[-30^{\circ},30^{\circ}]$, 2(b) scale on $x$-coordinate with scale factor $1.1^{\beta}$ where $\beta\in[-5,5]$, 2(c) adding noise with $\sigma/\sigma_{0}=[0,0.1]$, and 2(d) adding $n_{l}$ outliers where the outlier ratio$=n_{l}/n_{1}$.
@@ 4.3 CMU House and Hotel Dataset
Figure 4: The Cars and Motorbikes datasets with CURSOR and state-of-the-art hypergraph matching algorithms. 4(a) The number of non-zero compatibilities with ANN-based methods and CURSOR. The matching accuracy on the 4(b) Motorbikes and 4(c) Cars datasets.
@@ 6 PRL-based Method Derivation
Figure 6: The detailed hyperedge correspondences between one sampled hyperedge from the source (yellow triangle in 6(a)) and target images with ANN (6(b) and 6(c)) and CURSOR 6(d). The white triangles in the target images denote all the hyperedges compared with the source hyperedge, and the blue triangles are the hyperedges with the highest compatibilities ($r_{1}$ for ANN and $r$ for CURSOR). The green dashed triangles represent the matched hyperedges, and the red one represents the mismatch.
@@
We provide a more detailed analysis of the experiment results from Sec. 4.3 to demonstrate the superiority of CURSOR, as is shown in Fig. 6. Traditional ANN-based methods compute the compatibilities between the sampled source hyperedges (yellow triangle in Fig. 6(a)) and all the target ones (fully connected white triangles in Figs. 6(b)-6(c)). With the intermediate second-order result, CURSOR computes fewer compatibilities, as is shown in Fig. 6(d). To find the corresponding hyperedge (green dashed triangles in Fig. 6), a large amount (the hyperparameter $r_{1}=900$ in Fig. 6(b)) of the highest compatibilities (blue triangles in Figs. 6(b)-6(d)) should be selected for ANN. If we decrease $r_{1}$ to 50, some correct ones will be missed (red triangle in Fig. 6(c)). CURSOR can effectively find the hyperedges with the 25 highest compatibilities (green triangle in Fig. 6(d)). Compared to the traditional tensor generation methods, CURSOR can effectively increase the matching performance with less computational complexity.
@@ 8.1 CUR-based Pairwise Matching
Figure 7: 7(a) CUR-based second-order graph matching accuracy with various $c$. 7(b) The average accuracy using CURSOR with CUR-based pairwise matching result and CURSOR with randomly sampled indices (CURSOR-R).
@@ 8.2 CURSOR vs ANN-based Tensor Generation
Figure 8: Results on 100-vs-110 synthetic dataset comparing CURSOR with ANN, $\alpha=0.2$ for CURSOR. 8(a) The average matching accuracy using CURSOR and ANN with different $\alpha$. 8(b) The average iterations for convergence using CURSOR and ANN with different $\alpha$. 8(c) The decay $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}$ per iteration with different $\alpha$ using CURSOR and ANN. The noise level $\sigma=0.1$.
@@ 9.1 Parameter in Second-order Graph Matching
Figure 9: 9(a) The average hit rate over increasing $k$ on a 100-vs-110 synthetic dataset with $\sigma=0.02$ from the CUR-based second-order graph matching results. 9(b) The average accuracy on the 100-vs-110 synthetic dataset with $\sigma\in[0,0.2]$ and $c\in[10,100]$.
@@ 9.2 Parameters During Tensor Generation
Figure 10: Results on 100-vs-110 synthetic dataset with $\sigma\in[0,0.2]$ using CURSOR. 10(a) The matching accuracy with increasing $k$. 10(b) The matching accuracy with increasing $r$.
Figure 11: Results on 100-vs-110 synthetic dataset with PRL-based matching algorithm. 11(a) The average matching accuracy using CURSOR with different $\alpha$. 11(b) The average iterations for convergence using CURSOR with different $\alpha$. 11(c) The decay $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}$ per iteration with different $\alpha$ settings using CURSOR. The noise level $\sigma=0.1$.
@@ 9.3 Parameters of PRL-based Matching Algorithm
To study the sensitivity of $\alpha$, the matching accuracy and convergence speed using CURSOR were further analyzed. During the experiment, $c=100$, $k=10$, and $r$ was set as $100$. The stopping criteria of the PRL-based algorithm was assigned as $\|\mathbf{x}^{(k+1)}-\mathbf{x}^{(k)}\|_{2}\leq 10^{-8}$. The matching accuracy with various $\alpha$ is shown in Fig. 11(a). With a reliable compatibility tensor, the PRL-based matching algorithm achieved almost the same performance regardless of $\alpha$. We further analyzed their convergence speeds, as shown in Figs. 11(b) and 11(c). As the noise level increased, more iterations were required to satisfy the stopping criteria, and a lower $\alpha$ achieved faster convergence. As discussed in the previous sections, the parameter $\alpha$ is a balanced factor between first and third-order compatibilities. Although the method converged faster with $\alpha=0$, the first-order compatibilities stabilized the matching process and increased the matching performance under some extreme circumstances.